Following are the weights of 32 college students (in kg).
56 73 75 49 82 68 72 91 93 78 63 88 82 75 84 73

        67 68 79 91 82 76 80 67 79 80 91 86 90 67 82 75

Construct a 95% confidence interval estimate of the population standard deviation. Round your final answer to one decimal place, and write in interval notation.
Answer #1
56
73
75
49
82
68
72
91
93
78
63
88
82
75
84
73
67
68
79
91
82
76
80
67
79
80
91
86
90
67
82
75

\\ n=32\\\\ Mean(\bar{x})=76.94\\\\ Standard~Deviation(s)=10.386\\\\ degree~of~freedom=n-1=32-1=31\\\\ Critical~Value(t_{\alpha/2})=2.04\\\\ Confidence~Interval:\bar{x} \pm t_{\alpha/2}*s/\sqrt{n}\\\\ =76.94 \pm 2.04*10.386/\sqrt{32}\\\\ =76.94 \pm 3.745\\\\ =(73.195, 80.685)\\\\ =(73.2,80.7)